OpenStudy (anonymous):

find the critical numbers and the open intervals on which the function is increasing or decreasing...

OpenStudy (anonymous):

\[f(x)=x \sqrt{x+1}\]

OpenStudy (anonymous):

google it

OpenStudy (anonymous):

can you differentiate f(x) ?

OpenStudy (anonymous):

you'll need the derivative of f(x) and to find the critical numbers, solve f'(x)=0.

OpenStudy (anonymous):

do you have the derivative?

OpenStudy (anonymous):

yes i got\[\frac{ 1 }{ 2 }x (x+1)^{-\frac{ 1 }{ 2 }}\] is that correct? @byteme

OpenStudy (anonymous):

no... that's not right... you'll need to use the product rule/chain rule....

OpenStudy (anonymous):

\(\large f(x)=x\sqrt{x+1}=x(x+1)^{\frac{1}{2}} \) \(\large f'(x)=[x]' \cdot (x+1)^{\frac{1}{2}}+x\cdot [(x+1)^{\frac{1}{2}}]' \) = \(\large (x+1)^{\frac{1}{2}} + x \cdot \frac{1}{2}(x+1)^{-\frac{1}{2}} \cdot [x]' \) = \(\large (x+1)^{\frac{1}{2}} + x \cdot \frac{1}{2}(x+1)^{-\frac{1}{2}} \) = \(\large (x+1)^{-\frac{1}{2}}((x+1)^1 + \frac{x}{2}) \) = \(\large \frac{\frac{3x}{2}+1}{(x+1)^{\frac{1}{2}}} \)

OpenStudy (anonymous):

so \(\large f'(x)=0 \Rightarrow x=-\frac{3}{2} \)